# Large Language Models Understand Layout

### v1

@@ 5 Conclusion
This study extensively investigates the potential of LLMs in text layout understanding by constructing the TextLayoutQA dataset for in-depth research. Experiments utilizing various LLMs demonstrate that, compared to text without layout, the performance of LLMs on datasets with text layout improves by 8$\sim$25%, confirming their potential in text alignment, layout, and orientation understanding. The additional experiments show that during the pre-training phase, the base models already possess preliminary text layout understanding capabilities, which are further enhanced during instruction-tuning. Through ablation experiments with diverse instruction-tuning datasets, we find that training data is crucial for LLMs to acquire text layout understanding, particularly datasets containing text layouts such as codes. In addition, text layout understanding can be enhanced by low-cost auto-generated data approached by a novel text game. Subsequently, leveraging the text layout understanding capabilities of LLMs, we propose an approach named TextLayoutParser to address text-rich VQA problems, achieving decent performance improvements on the XfundQA, FetaQA, and DocVQA datasets.
@@ 6 Acknowledgments
This research was supported by "Pioneer" and "Leading Goose" R&D Program of Zhejiang (No. 2024C01020).
@@ Appendix A Examples of LLMs Text Layout Understanding Capability
During the early exploration of GPT-3.5-Turbo’s QA ability, its remarkable capability to comprehend text alignment, layout, and orientation was discovered. Figure 6 shows some examples of the exploration.
@@
Figure 6: Showcases of GPT-3.5-Turbo’s text layout understanding capability.
@@ Appendix B Corpora with Layout Information on Github and StackExchange
By searching for data within Pile that potentially contains text layout information, we discover considerable relevant data from sources like GitHub and StackExchange. Figure 7 shows some examples.
@@
Figure 7: An example of data with layout information on Github and StackExchange.
Table 13: Test results, measured by ROUGE-L and recall, of using different types of instructions to tune LLMs. The "Instructions" column specifies the different instruction-tuning datasets. "Origin" indicates no instruction-tuning performed. "Base", "Code", "Table", and "Generate" correspond to instruction-base, instruction-table, instruction-code, and instruction-generated datasets, respectively. The "Tasks" row specifies the different subsets within the test set, with "Others" encompasses results excluding "Code", "Table", and "Generate" subsets.
@@
Appendix C Test results of using different types of instructions to tune LLMs
The test results of using different types of instructions to tune LLMs are presented in Table 13. Given LLMs often produce long responses that don’t align with the ground truth of the table subset, recall yields more reasonable results than ROUGE-L. It can be observed that, compared to the chat model, the code capability significantly decreases after tuning on instruction-base, but it substantially recovers after tuning on the instruction-code, except for Baichuan2-7B. In contrast to the model tuned on the instruction-base, table capability gains a considerable improvement after tuning on the instruction-table, while layout capability obtains a remarkable improvement following instruction-generated tuning.
Appendix D Method of textLayoutParser
@@
Appendix E Prompt Designs for Difference Datasets
Figure 8 illustrates the prompt designs for different datasets. (a) display one-shot prompting for TextLayoutQA. (b)$\sim$(d) display zero-shot prompting for XfundQA, DocVQA, and FeTaQA, respectively. (e) illustrates the 3-shot prompting for rephrasing answers in the DocVQA dataset. The instructions remain consistent across all LLMs except for the Llama2 series, as depicted in (f). Regarding LLM parameter settings, we utilize a temperature of 0.1, maximum output length of 512, top p of 0.85, and repetition penalty of 1.05.
@@ Appendix E Prompt Designs for Difference Datasets
Figure 8: Prompt designs for different datasets.
Appendix F Examples of different table encoding methods
Figure 9 shows examples of different table encoding methods. The widely used table encoding methods include: arranging data in array format (Array), using unique identifiers to distinguish between headers and rows (Linear), and formatting each element as a column-row-value triplet to form a list (Triple).
@@ Appendix F Examples of different table encoding methods
Figure 9: Different table encoding methods: "Array," which transforms the original array table data into string format; "Linear," which employs distinct identifiers to differentiate headers and rows; "Triplet," which formats each element as a col-row-value triplet to create a list; and "Ours," which utilizes spaces and line breaks to align and separate elements within the table.

### v3

@@ 5 Conclusion
This study extensively investigates the potential of LLMs in text layout understanding by constructing the TextLayoutQA dataset for in-depth research. Experiments utilizing various LLMs demonstrate that, compared to text without layout, the performance of LLMs on datasets with text layout improves by 8$\sim$33%, confirming their potential in text alignment, layout, and orientation understanding. The additional experiments show that during the pre-training phase, the base models already possess preliminary text layout understanding capabilities, which are further enhanced during instruction tuning. Through ablation experiments with diverse instruction-tuning datasets, we find that training data is crucial for LLMs to acquire text layout understanding, particularly datasets containing text layouts such as codes. In addition, text layout understanding can be enhanced by low-cost auto-generated data approached by a novel text game. Subsequently, leveraging the text layout understanding capabilities of LLMs, we propose an approach named TextLayoutParser to address text-rich VQA problems, achieving decent performance improvements on the XfundQA, FetaQA, and DocVQA datasets.
@@ 6 Acknowledgments
This research was supported by "Pioneer" and "Leading Goose" R&D Program of Zhejiang (grant no. 2024C01020).
@@ Appendix A Examples of LLMs Text Layout Understanding Capability
During the early exploration of GPT-3.5-Turbo’s QA ability, its remarkable capability to comprehend text alignment, layout, and orientation was discovered. Figure A1 shows some examples of the exploration.
@@
Figure A1: Showcases of GPT-3.5-Turbo’s text layout understanding capability.
@@ Appendix B Corpora with Layout Information on Github and StackExchange
By searching for data within Pile that potentially contains text layout information, we discover considerable relevant data from sources like GitHub and StackExchange. Figure A2 shows some examples.
@@
Figure A2: An example of data with layout information on Github and StackExchange.
Appendix C Parameters of the instruction-tuning
Table A1: Key parameters of the instruction-tuning
parameter value
num_train_epochs 5
learning_rate 1e-4
max_seq_length 1024
lr_scheduler_type constant_with_warmup
warmup_steps 1000
lora_rank 64
lora_alpha 16
lora_dropout 0.05
gradient_checkpointing true
optim adamw_torch
fp16 true
weight_decay 0
max_grad_norm 0.3
Table A2: Test results, measured by ROUGE-L and recall, of using different types of instructions to tune LLMs. The "Instructions" column specifies the different instruction-tuning datasets. "Origin" indicates no instruction-tuning performed. "Base", "Code", "Table", and "Generate" correspond to instruction-base, instruction-table, instruction-code, and instruction-generated datasets, respectively. The "Tasks" row specifies the different subsets within the test set, with "Others" encompasses results excluding "Code", "Table", and "Generate" subsets.
@@ Appendix C Parameters of the instruction-tuning
Appendix D Test results of using different types of instructions to tune LLMs
The test results of using different types of instructions to tune LLMs are presented in Table A2. Given LLMs often produce long responses that don’t align with the ground truth of the table subset, recall yields more reasonable results than ROUGE-L. It can be observed that, compared to the chat model, the code capability significantly decreases after tuning on instruction-base, but it substantially recovers after tuning on the instruction-code, except for Baichuan2-7B. In contrast to the model tuned on the instruction-base, table capability gains a considerable improvement after tuning on the instruction-table, while layout capability obtains a remarkable improvement following instruction-generated tuning.
Appendix E Method of textLayoutParser
@@
Appendix F Prompt Designs for Difference Datasets
Figure A3 illustrates the prompt designs for different datasets. (a) display one-shot prompting for TextLayoutQA. (b)$\sim$(d) display zero-shot prompting for XfundQA, DocVQA, and FeTaQA, respectively. (e) illustrates the 3-shot prompting for rephrasing answers in the DocVQA dataset. The instructions remain consistent across all LLMs except for the Llama2 series, as depicted in (f). Regarding LLM parameter settings, we utilize a temperature of 0.1, maximum output length of 512, top p of 0.85, and repetition penalty of 1.05.
@@ Appendix F Prompt Designs for Difference Datasets
Figure A3: Prompt designs for different datasets.
Appendix G Examples of different table encoding methods
Figure A4 shows examples of different table encoding methods. The widely used table encoding methods include: arranging data in array format (Array), using unique identifiers to distinguish between headers and rows (Linear), and formatting each element as a column-row-value triplet to form a list (Triple).
@@ Appendix G Examples of different table encoding methods
Figure A4: Different table encoding methods: "Array," which transforms the original array table data into string format; "Linear," which employs distinct identifiers to differentiate headers and rows; "Triplet," which formats each element as a col-row-value triplet to create a list; and "Ours," which utilizes spaces and line breaks to align and separate elements within the table.